OpenStudy (anonymous):

2(5-3x)=1/16 solve (just give me the steps for starters, I think I can get it

OpenStudy (anonymous):

I tried multiplying the two into the parenthesis, moving the 10 over and dividing out the -6 from both sides, but that didn't seem to work. got 53/32

OpenStudy (anonymous):

i hope this is right..if not i apologize!! :o first you subtract 5-3 then you get 2 then multiply 2 and 2x you get 4x then you have 4x=1/16 divide 1/16...you get 16 then you have 4x=16 divide 16/4 and you come out with 4=x

OpenStudy (anonymous):

2(5-3x)=1/16 2(2*4)=1/16 2(8)=1/16 16=1/16 16=16

OpenStudy (anonymous):

hmm, well, unfortunately that answer is not one of the options. it's either -3, 1/8, 3 or 8

OpenStudy (anonymous):

im afraid you've either typed in the question in correctly or those arent the right options for answers... 2 (5 - 3x) = 1 / 16 10 - 6x = 1 /16 16 (10 - 6x) = 1 160 - 96 x = 1 -96 x = 1 - 160 96x = 159 x = 159 / 96 x = 1.65625

OpenStudy (anonymous):

@ChristopherAdams yea, thats the same thing I did, and got the same. I guess I'll have to email my teacher about it. thanks